OpenStudy (anonymous):

The sun roof in Troy’s house is shaped like a rectangle. He increases the length of the sides of the sun roof to twice the size of the existing ones. How will this change affect the perimeter of the sun roof?

OpenStudy (jlastino):

The perimeter of a rectangle = 2(l +w ) When you increase the size of the length and width it would be 2(2l+2w) Lets assign values to l and w where l>w l=4 l=2 The original roof would have a perimeter of 2(4+2)= 12 Doubling the l and w would give us 2(4(2)+(2)2)= 2(8+4) =24 From this we can conclude that the perimeter doubled

OpenStudy (anonymous):

thanks man